OpenStudy (xapproachesinfinity):

what method can i use the factor this polynomial X^4-12x^3+59x^2-138x+130 this one has no real zeros, the zeros are in complex system.

OpenStudy (xapproachesinfinity):

@zepdrix

OpenStudy (xapproachesinfinity):

I know the answer already, I want to know how to get it

zepdrix (zepdrix):

Hmm not sure :C thinking....

OpenStudy (xapproachesinfinity):

take your time! I got this in exam. there was multiple choices i went ahead a looked at the ones that make sense and FOIL ed them, so I can get the original one haha. out of 4 choices a selected two that make sense. I got the right answer this way, but I'm unsatisfied haha

OpenStudy (xapproachesinfinity):

I was playing probability gamble lol

zepdrix (zepdrix):

If you toss it into Wolfram, they factor it into,\[\Large\rm \left[(x-6)x+10\right]\left[(x-6)x+13\right]=0\]And then further into:\[\Large\rm \left[x^2-6x+10\right]\left[x^2-6x+13\right]=0\]And from there it's easy peasy finding the complex roots. I'm not sure how they found the x-6 factors though. wio or hartten might have some crazy method for doing this.

OpenStudy (xapproachesinfinity):

Oh! I was playing with something like this I did get x-6 thing in one side but i was not sure where it was leading so I give up. Thanks! let's ask these guys for more tips so I clear this for ever lol

OpenStudy (xapproachesinfinity):

@wio, @hartten

OpenStudy (xapproachesinfinity):

I will check what i did again and see

zepdrix (zepdrix):

offline? darn :c

OpenStudy (xapproachesinfinity):

No worries, and thanks a lot.

OpenStudy (xapproachesinfinity):

@ganeshie8 some leads here!

OpenStudy (anonymous):

:o I also want to know how to get (x-6). Looks like a difficult but fun problem

OpenStudy (xapproachesinfinity):

I almost got something like that but time ran off it was an exam haha I was like darn it, let's just use some other fake ways and get the result haha So I foiled two options that make sense and get the answer, others were using TI to get that Factorization form zepdrix wrote

OpenStudy (xapproachesinfinity):

i remember i split 59 to 36 and 23 so i can complete the square with x^4-12x^3 part

OpenStudy (xapproachesinfinity):

i got x^2(x-6)^2

OpenStudy (xapproachesinfinity):

i need to do something similar with the other stuff haha, but time was going against me

OpenStudy (xapproachesinfinity):

needed*

OpenStudy (xapproachesinfinity):

@mathmale any idea to easily get this?

OpenStudy (rational):

familiar with below precalc goodies ? 1) rational root theorem 2) descartes rule of signs 3) synthetic division 4) factor/reminder theorem

OpenStudy (rational):

nvm, none of them gona work >.<

OpenStudy (xapproachesinfinity):

i'm familiar with those but they don't work here. if i could find one complex zero this solved done! but I could use any of those theorem. however I might be wrong, so i would like you to try it and see

OpenStudy (xapproachesinfinity):

@ikram002p what can say?

OpenStudy (rational):

so basically you want some stable method to factor it into two quadratics ?

OpenStudy (xapproachesinfinity):

Correct!

OpenStudy (ikram002p):

since it has no real zeross , is it ok to give two quadratic combination ?

OpenStudy (xapproachesinfinity):

Yes! i needed a stable way to derive those quadratics

OpenStudy (rational):

finding the quadratic factors is the tricky thing

OpenStudy (ikram002p):

mmmm ic :D it might be more than one solution

OpenStudy (xapproachesinfinity):

it is! it involves may steps as i said i was working toward that but i give up haha

OpenStudy (rational):

im thinking something like this may work : \[x^4-12x^3+59x^2-138x+130 = (x^2+ax+b)(x^2+cx+d)\]

OpenStudy (xapproachesinfinity):

ikram you mean the zeros? they are 4!

OpenStudy (ikram002p):

but , tbh that wont be called factorize any more :o its only give combinationns xD

OpenStudy (rational):

compare the coefficients both sides and solve FOUR freaking equations !

OpenStudy (ikram002p):

then FOIL wow that might work ! rashhhhhhhh

OpenStudy (xapproachesinfinity):

actually u are right ikram the question asked linear combinations

OpenStudy (xapproachesinfinity):

rational, that sound good let me check it!

OpenStudy (xapproachesinfinity):

ikram i foiled hhh that's how i got the right answer i didn't have time for it hhh

OpenStudy (xapproachesinfinity):

oh you're talking about what ratio said lol

OpenStudy (xapproachesinfinity):

okay one second to see if that works

OpenStudy (ikram002p):

rashhh only Foil well not bad having 4 equation :3

OpenStudy (xapproachesinfinity):

wow so many equation haha

OpenStudy (rational):

they don't look pleasant haha! \(-12 = c+d\) \(59 = d + ac + b\) \(-138 = ad +bc\) \(130 = bd\)

OpenStudy (xapproachesinfinity):

a+c=-12 no?

OpenStudy (rational):

yes :) but lets give it up, they don't look solvable \(-12 = c+a\) \(59 = d + ac + b\) \(-138 = ad +bc\) \(130 = bd\)

OpenStudy (xapproachesinfinity):

it's too much haha

OpenStudy (ikram002p):

you give up, wth ! its only FOUR equations

OpenStudy (rational):

the original question was only one equation :P

OpenStudy (xapproachesinfinity):

I don't like that equation haha

OpenStudy (xapproachesinfinity):

but the way sounded nice haha, thanks ratio

OpenStudy (ikram002p):

wolfram ?

OpenStudy (xapproachesinfinity):

okay how do we do it? actually i have never heard of wolfram haha, may be i used but don't know what it is lol my background is different

OpenStudy (ikram002p):

OMG! hehe it did as what i thought xD but i thought that was silly so didnt post hehe

OpenStudy (xapproachesinfinity):

that's cheating lol haha

OpenStudy (rational):

what it did ?

OpenStudy (ikram002p):

i start from bd=130 and give possible values for b and d

OpenStudy (ikram002p):

wolfram fixed b

OpenStudy (rational):

yeah that would be guess and check again, not systematic

OpenStudy (ikram002p):

hehe

OpenStudy (xapproachesinfinity):

I actually have no idea how that works?

OpenStudy (ikram002p):

mmm forget it lol why it would be important to give combination anyway :P

OpenStudy (xapproachesinfinity):

haha, the question asked to do so, and i got 4 options

OpenStudy (ikram002p):

what are they ?

OpenStudy (xapproachesinfinity):

they are too much to write haha 4 linear combinations all of them i foiled two of to get the answer easier than trying to get that first factorization haha

OpenStudy (xapproachesinfinity):

prof is expecting us to use the Texas Instrument

OpenStudy (ikram002p):

what are the zeros atleast?

OpenStudy (ikram002p):

:o Texas

OpenStudy (rational):

hey @ikram002p actually your factoring is the best method here as 130 has very few factors and we can find the values of a,b,c,d very easily...

OpenStudy (xapproachesinfinity):

the zeros are x=3-i, 3+i, 3-2i, 3+2i the question is saking factor as a product of linear factors

OpenStudy (rational):

b and d can be one of : {2,5, 13}

OpenStudy (ikram002p):

ic:o then it would be four possible equation ,wolfram gives 2

OpenStudy (xapproachesinfinity):

other students used TI while i was struggling haha

OpenStudy (xapproachesinfinity):

it should be 4 possible zeros since it is a 4th degree poly

OpenStudy (ikram002p):

idk what is texas or TI

OpenStudy (ikram002p):

4 possible linear combination ,this is what i ment

OpenStudy (xapproachesinfinity):

Graphing calculator

OpenStudy (xapproachesinfinity):

i don't have one hehe

OpenStudy (ikram002p):

ohh then wolfram is not cheating :P since u suppose to solve it with calculator

OpenStudy (xapproachesinfinity):